OpenStudy (anonymous):

Under ideal conditions a certain bacteria population is known to double every 2 hours. Suppose there are initially 200 bacteria. What is the size of the population after t hours?

OpenStudy (anonymous):

@Hero

OpenStudy (wolf1728):

Actually, after 6 hours population = 1,600

OpenStudy (wolf1728):

Yes, I was working on a formula to :-( Seems we lost Hero?

OpenStudy (anonymous):

i know the 6 hours already but i need the formula

OpenStudy (anonymous):

it says its wrong

OpenStudy (wolf1728):

No because after 2 hours population = 200 * 2² = 400? But problem states population = 200

OpenStudy (wolf1728):

2 Hours Population = 200 * 2^0 4 Hours Population = 200 * 2^1 6 Hours Population = 200 * 2^2 8 Hours Population = 200 * 2^3

OpenStudy (wolf1728):

Population = 200 * 2^((t-2)/2) ???

OpenStudy (anonymous):

no thats not it either

OpenStudy (wolf1728):

You're right muzz I thought this was going to be easy.

OpenStudy (wolf1728):

After 4 hours (4-2/2) = 2 Raising that to the power of 2 is 4 times 200 = 800 no

OpenStudy (wolf1728):

2 4 6 8 0 1 2 3 Population = 200 * 2^((t-2)/t) ???

OpenStudy (wolf1728):

No that doesn't work for 4

hero (hero):

x|y 0|200 <200*2^0> 2|400 <200*2^1> 4|800 <200*2^2> 6|1600 <200*2^3>

hero (hero):

200^(t/2)

OpenStudy (wolf1728):

2 4 6 8 0 1 2 3 No it can't be 200^(t/2)

OpenStudy (anonymous):

thanks @Hero

OpenStudy (anonymous):

Estimate the size of the population after 19 hours. i got 154,000. whats wrong

OpenStudy (wolf1728):

muzz you think 200^(t/2) is the answer?

OpenStudy (wolf1728):

I thhink that formula is wrong

hero (hero):

It's right. I'm pretty sure @muzzammil.raza inputs his answers into some computer program that verifies the correct answer.

OpenStudy (anonymous):

no the homework is online

OpenStudy (anonymous):

it tells u if u r right or wrong. u have a week to finish it

OpenStudy (anonymous):

Estimate the size of the population after 19 hours. i got 154,000. whats wrong

OpenStudy (wolf1728):

200 * 2^0 = 2 hours 200 * 2^1 = 4 hours 200 * 2^2 = 6 hours 200 * 2^3 = 8 hours Just have to figure out a formula for this.

hero (hero):

That online homework that tells you if you are right or wrong is the computer program I'm referring to @muzzammil.raza

hero (hero):

We already figured out the formula. x|y 0|200 <200*2^0> 2|400 <200*2^1> 4|800 <200*2^2> 6|1600 <200*2^3> 8|3200 < 200*2^4>

hero (hero):

@wolf1728, his program verified that 200^(t/2) is correct.

OpenStudy (wolf1728):

200 * 2^(19/2) 200 * 2^9.5 Ppopulation after 19 hours = 200 * 724.077343935 = 144,815.468787005

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

200 ^ (t/2) is correct?? then after 8 hours population is 200^4?? =1,600,000,000 After 8 hours it is 1,600

hero (hero):

yeah, it was 200*2^(t/2) sorry

hero (hero):

This is clearly my biggest screw up of the day.

OpenStudy (wolf1728):

Yeah Hero I'm surprised at myself too. I thought I'd type a formula in a few seconds. LOL

hero (hero):

Well, as long as we figured it out. I fried a few brain cells in the process for sure.

OpenStudy (wolf1728):

Well it is done !!! :-)

OpenStudy (wolf1728):

muzz is back - I sent you a message

OpenStudy (wolf1728):

And I just got the answer - okay !